Find an equation of the line satisfying the conditions. Write the equation in slope-intercept from. Through (-6,7); parallel to 3x+7y=3
Answer by fractalier(6550)

The slope of 3x+7y=3 is -3/7, found by m = -A/B
Since we want a line parallel, slopes are equal...
Now use the slope-intercept form
y - y1 = m(x - x1)
y - 7 = (-3/7)(x + 6)
y - 7 = (-3/7)x - 18/7
y = (-3/7)x + 31/7 (since 7 = 49/7)